The Trial

William Astle February 5, 2004

Notes

This work is copyright (c) 2004 by William Astle.

This work is licensed under the Creative Commons Attribution-NoDerivs-NonCommercial License. To view a copy of this license, visit http://creativecommons.org/licenses/by-nd-nc/1.0/ or send a letter to Creative Commons, 559 Nathan Abbott Way, Stanford, California 94305, USA.

I am also greatly interested in any feedback anyone may have. You can send such feedback by email to lost@l-w.net. You may also be able to find more works by me at my web site at http://lost.l-w.net/ so please stop by.

The Trial

”This court will come to order!” the magistrate boomed. A hush fell on the great chamber as everyone looked to the front of the room with anticipation. The magistrate continued in a lower tone of voice, ”We are gathered today to determine the fate of one individual whose only crime was to be in the wrong place at the wrong time. We will also be determining what action will be taken against the two of our own for precipitating the events with which we are concerned today.”

As the magistrate finished his introduction, two men were escorted into the chamber and directed to seats on a raised dais near the front of the room. Immediately beyond the dais, the council bench loomed over it. One man was relatively young with dark hair and a short beard. His expression was grim but the gaze he sept over the gathered crowd was clear and unwavering. He had a name tag which read, ”Sola”. The other was older, with a few traces of gray showing in his red hair. His eyes revealed the full force of the rage which had led to his appearance here. Yet only his eyes showed any emotion as he entered the room and sat down. He also sported a name tag. His read, ”Regat”.

When the two men were seated, the magistrate continued, ”The Wizard Sola and the Sorcerer Regat stand accused of conducting a duel in a location which threatened the safety of noncombatants. While these charges would be grave under ordinary circumstances, the consequences of this particular duel involved an individual from a planet not part of the Confederation. Indeed, the very course of the history of the planet may have been affected by these actions. Neither combatant disputes the occurrence of the duel and thus we need not concern ourselves with the guilt or innocence of the combatants. We will, instead, concern ourselves with determining exactly what happened and what action must be taken as a result.

”In light of the severity of this incident, the council has approved the use of the past viewing technique to obtain an objective view of events.” There was a collective intake of breath at this pronouncement. Even the combatants themselves showed some level of surprise. ”It was determined that because the fate of an entire planet may be at stake, the enormous energy expense for this procedure is warranted. We will start the viewing immediately before the start of the duel.”

At this point, a loud mechanical sound permeated the council chamber. Slowly, a large projection arrangement lowered from the ceiling and stopped about three metres from the floor with a loud clang. Then, anticlimactically, dead silence greeted the sudden appearance of the three dimensional image immediately below the projector. The projection scheme itself was purely technological; after all, there was no need to expend magical energy to do something so mundane as making an image. Indeed, the projection device was used for more than just past viewings; it served to project any images necessary in the council chamber. The real magical component was the past viewing itself which simply fed its results into the projector.

The scene before the assembly was that of a small laboratory, littered with the usual paraphernalia that one found in such a place. That is, benches, test tubes, rube goldberg like apparatuses that one would be hard pressed to identify a purpose for, and so on. Standing just within the door to the laboratory were the two defendants. They were having a conversation, the contents of which were clearly audible to the audience.

”Why, precisely, should magic and science be regarded as separate disciplines?” Wizard Sola was saying. ”I mean, they both exist in the universe. There must be scientific principles which guide both.”

”Why must there be? All through the ages, magic has been singularly inexplicable to the scientists. They have yet to even come up with an adequate theory to describe what we can see about the universe, let alone magic,” Regat replied.

”Since when did not knowing the answer mean the answer did not exist?” Regat had no answer for this. The conversation continued for a few minutes

before returning to the subject with Sola saying, ”Well, I intend to organize other like-minded individuals, scientists and magic users alike, to research this idea. We will demonstrate that magic is not inexplicable to the sciences.”

”Why must you continue with this ridiculousness? You’ll only bring yourself frustration.”

”At least I am willing to question what has always been! Neither magic nor science can possibly advance if nobody is willing to question what is!” A note of frustration mixed with anger crept into Sola’s voice as he said this.

Responding to the tone of Sola’s voice, Regat replied, ”Are you insinuating that those of us who have spent our lives mastering our arts are not of the same calibre as such impetuous young ones as yourself?”

The two of them traded heated remarks for several minutes, the tone getting steadily more serious. Then, eventually, Sola said calmly, ”Neither one of us is served by this arguing. Let’s agree to disagree before something untoward happens.”

Regat, however, would have none of it and simply roared. As he roared, he made several expansive gestures with his arms and Sola was hurled across the lab, breaking several apparatuses on the way. Sola landed hard and had the breath knocked out of him. As he struggled to his feet and tried to catch his breath, a lance of flame streaked toward him from Regat’s finger. Forced to defend himself, Sola raised his arm in a gesture of warding and the fire was deflected, demolishing a nearby workbench. By this time, he had recovered his breath enough to say, ”Regat, have you lost your mind!? This is no setting for a duel!”

Regat growled and attacked again. Realizing that he had no choice but to fight back, Sola employed all his knowledge of magic to defend himself. He also called upon every bit of literature he had ever read, searching for a way to beat Regat without killing him and without endangering anyone. As the battle raged on, minutes seeming to be hours, Sola was becoming more and more desperate. He was no match for an opponent enraged beyond reason. Indeed, to the audience of this spectacle, it seemed clear that Regat had succumbed to a berserker rage.

Suddenly, and surprisingly, Sola simply stopped. It seemed he had ceased to perform any magic at all. This only lasted for a few seconds as Regat prepared another attack but it seemed that Sola had lost his mind. Then it became apparent that Sola had been up to something. Just as Regat unleashed his next attack, Sola spread both his arms wide. There was a brilliant flash which left most of the audience temporarily blinded and the two combatants were gone. They reappeared moments later with Regat sprawled on the floor with Sola staggering with exhaustion and looking very disappointed. The two duellists were shortly joined by appropriate authorities and taken into custody as the projection faded.

The magistrate then addressed the gathering, ”The point at which the duel disappeared from the laboratory is the key point with which we are interested today. The events as displayed clearly indicate that Regat was the aggressor and Sola merely defended himself. Were that the only consideration, this proceeding would be moot. We will now show the sequence of events from the point of the disappearance until the reappearance in the laboratory. This sequence is somewhat confusing as the background will be changing fairly rapidly as the duel ensues.”

The projector sprang to life again. This time, the image started immediately before the blinding flash but instead of staying with the laboratory, the image stayed with the combatants. The background changed abruptly from the battered lab to an open field with a blue sky overhead. Sola and Regat maintained their relative positions during this transition. Sola continued with his arms spread while Regat threw fireballs at him. A few seconds later, the background shifted abruptly again, this time to an infinite blackness with no means of orientation. Both combatants appeared discomfitted by this change, as though they couldn’t breathe.

The rapid fire change of scenery continued for about a minute, with settings as varied as open fields, underwater vistas, forests, and magnificent cities. As this continued, Sola looked more and more drained while Regat continued to hurl attack after attack at him. Regat was looking more and more determined as the battle raged but the determination was tinged increasingly with disbelief. None of his attacks since the scene shifting had started had come anywhere near close to scoring on Sola.

Then the shifting background stabilized for a moment. The two duellists had appeared in mid air over a mountain range. A large river was visible in the distance. Both parties immediately began levitating to prevent their own deaths. Then the battle took on a new dimension. Regat, at this point realizing he was overmatched by Sola, even though he could not comprehend why, fled with Sola streaking after him. A net of bright white reached forward from Sola’s hands to envelop Regat.

The action paused at this point as the magistrate said, ”Please pay attention to the ground level from this point forward.” The action resumed.

As the streaking combatants raced through the skies of an unknown world, what appeared to be a road came into view on the ground. On a flat between mountain ranges was a large cleared area which appeared to be built up similar to the road. There was a corner immediately to one end of the cleared area. Halfway around the corner was some sort of vehicle which was glowing brightly in an electric blue light. The vehicle was not moving at all although the sharp eyed among the crowd could see the occupant moving. All this was only visible for about ten seconds as the battle in the sky streaked past even as it drew to a close as the net of light enveloped Regat and the background changed back to the devastated laboratory and the projection faded.

”Before we view the events as seen by the occupant of the vehicle in the final location of the duel, we have a few questions to ask of the participants. Wizard Sola, would you please describe what you were trying to do immediately before the romp through various scenery?”

Sola’s voice was strong and unwavering as he said, ”I realized I didn’t stand a chance against what seemed to me to be a berserker rage. I felt that my only chance to come out of the situation alive was to put absolutely everything I had into a shield. Even then, I thought I had no chance of actual survival.”

”Do you have any idea what happened to cause the wild journey that followed?”

”As I released my energy into the shield, I felt Regat release his attack. That, in itself, would not have caused anything as wild as all that; he was simply using conventional magical attacks. However, at the exact instant that my shield and Regat’s attack converged, I felt another force - that’s the best I can do to describe it, even now - tugging at us and next thing I knew, the scenery around us was changing rapidly. In a few instances, our surroundings were inimical to oxygen breathing creatures. Fortunately, we did not stay long in any of those locations. As to where any of them were, I have no idea.”

”And you, Sorcerer Regat, have you any idea what might have happened to cause this journey?”

Regat, still looking somewhat stormy, replied with an even voice that betrayed no hint of anger, ”I also felt that third force and have no speculation as to what could possibly have caused it. I also concur that nothing either myself or Sola was doing at the time could have caused the location shifting. I can confirm that Sola was creating a shield - I felt it when my attack connected with it.”

”Did either one of you notice the vehicle on the road in the final location of the location shifting?”

Both defendants shook their heads. Sola then said, ”I did notice that the third force was strongest in that location. Perhaps there is a connection?” Regat looked at Sola with surprise in his expression and voiced his agreement with the assessment.

”Thank you for your candour. We will now view the incident from the point of view of the occupant of the vehicle.” Then the projector came to life once more, this time providing a view from within the vehicle. The view jumped all over as the occupant looked around at the various instruments that were used in operating the vehicle as well as the scenery and the road itself. Once the viewers became used to the point of view shifting around in a natural manner, it was easy to follow the action. The vehicle was making rapid progress along a gravel road. Then, suddenly, half way around a corner, the vehicle stopped. There was no corresponding lurch in the viewpoint to indicate a rapid stop. The viewpoint moved past all the various gauges and the scenery. It settled for about half a minute on what appeared to be a chronometer on the occupant’s wrist. The chronometer was showing no changes. (”The symbols on the right hand side of the display should be changing at a constant rate that would be visible even during this brief view,” the magistrate put in at this point.)

Eventually, the viewpoint moved upward. There, a brilliant streak of light crossed the field of view from the right to the left in just ten seconds or so. A very short time after this, the vehicle was in motion again as though it had never stopped and the viewpoint resumed it’s pre-interruption pattern until the vehicle pulled into what was clearly a parking lot now that it was seen from ground level rather than from the air (where it had appeared as a clearing). The vehicle was stopped and the viewpoint skewed as the viewer got out of the vehicle and took on a regular pattern which one could associate with walking. The view then faded as the projector was switched off again.

”It is not clear what happened to the occupant of that vehicle as a result of this incident. Indeed, it is not clear whether it was the duel which caused the occupant’s involvement or if the occupant somehow pulled the duel into his location. (And yes, we do know that the occupant of the vehicle is male.) We do know that there is no evidence of detrimental effects to this individual but there has been insufficient time for many possible effects to manifest. Does either of the defendants have anything to add to these proceedings before the council convenes to discuss the situation?”

Regat’s eyes had calmed almost completely over the course of the brief ride with the vehicle’s occupant. He and Sola shared a brief glance and Regat nodded imperceptibly. This action confused the audience significantly but apparently Sola knew exactly what Regat meant. Sola then said, ”It seems to me that something beyond the experiences of anyone currently practising magic has occurred. I believe my colleague here would agree that further study of the incident is warranted, whatever the outcome of this proceeding. I would personally like to volunteer to participate in the study although I cannot speak for my colleague.” Regat simply nodded again.

”Thank you for your contribution to this inquiry. Does anyone here have anything of relevance to add?” Nobody in the audience made any move to speak. Such a request was nearly unheard of although it was allowed by the procedures. Given the nature of this inquiry, it was not particularly surprising. Of course, the same thing that made the request unsurprising made the lack of response unsurprising as well; nobody had any idea what to make of the situation. ”In that case, the council will retire to discuss what has been learned today. The council will present its findings seven days hence.”

Seven days later, the audience in the council chamber was an order of magnitude larger and filled every conceivable cubic metre of the chamber and even a few inconceivable cubic metres. The defendants again occupied their chairs on the dais and the council was again arrayed at its bench at the front of the room. The magistrate called the court to order and then turned the floor over to one council member who had particularly white hair and a long beard. He shuffled forward, leaning heavily on his staff, to the podium that the magistrate had occupied previously (just in front of the centre of the council bench).

”After many long hours of discussion on this matter, the council has come to the following conclusion,” the old man said in a remarkably strong voice. ”The council has concluded that in the case of the duel, the Sorcerer Regat is fully at fault for the duel occurring. While this is a serious breech of protocol, it would not have affected anyone incapable of defending themselves had it remained in the laboratory until its conclusion. Given Regat’s exemplary record previously, we have opted to reprimand him sternly with a warning not to let his temper get him into another similar situation. We also find that Sola is entirely blameless in the matter of the duel.” Thus far, there were no surprises.

”The matter which caused us considerable debate, however, was twofold. What to do about the apparently innocent bystander and what caused the location shifting. Unfortunately, there is no way we can effectively figure this out after the fact without duplicating the event. The council debated long to reach this conclusion. We do, however, agree that significant study needs to be dedicated to this phenomenon. To that end, Wizard Sola and Sorcerer Regat are requested to form a team to research it.” There was a collective sense of surprise at this pronouncement. In response, the old man noted, ”There are no others even remotely qualified to direct this research, so do not take it as any kind of reflection of our opinion of your ability.”

The old man paused for a moment. Then he continued, ”During our deliberations, we spent some time identifying the location of that final scene in the duel. We have narrowed it down to a planet known as Earth whose inhabitants do not, as a whole, believe in magic nor do many of them possess any amount of magical ability. The ability is not totally unknown among them and paradoxically, those who do possess it usually possess a great deal of it. They are currently in the early electronics age of technological development but they are progressing extremely rapidly. Their political system is very volatile and they are a very violent group as a whole. The location on Earth where the incident occurred was isolated so fortunately very few people will even have noticed anything.

”We also did some scrying and discovered that the apparent bystander is one of those rare individuals who possess magical talent.” There was a collective intake of breath at that. Even the two defendants were visibly surprised. ”We suspect that the glow we saw around the vehicle was possibly an unconscious self defence mechanism in operation. It may be possible that when Sola was gathering strength for his shield that he managed to link with this individual somehow and this may have been the cause of much of the confusion. However, much study will be required to determine the exact mechanism that was at work.

”Because the Terran affected by all this already possessed magical ability, we see no reason to intervene in the affairs of Earth. However, we have undertaken to monitor events on Earth both because of this incident and the fact of the powerful magical ability that some of its inhabitants possess. The council has no further comment on this matter.”

The old man resumed his original seat among the council as the magistrate resumed the podium. While this was happening, Sola and Regat had a quiet conference and Sola stood up. The magistrate recognized Sola and then Sola said, ”My friend and I have decided that it is likely we will be able to carry out our research more effectively on Earth where we may observe these Terrans. We feel that understanding Terrans may help us understand what happened to us. To that end, we volunteer to head up not only the research into the phenomenon we experienced but also the observation of Earth.”

The council hastily conferred among themselves and then one by one nodded their assent. The magistrate then said, ”It appears the business before this gathering is complete. Our best wishes go with Wizard Sola and Sorcerer Regat. We are adjourned.”

There was much discussion among the populations of the various worlds of the Confederation for the weeks and months that followed. Had any of the Terrans a clue how much discussion they had managed to provoke out among the stars by simply existing, they would likely have been amused. Some, perhaps, might have been astounded and others thrilled, but mostly they would have been amused. After all, most of them knew, deep down, how insignificant their little world was in the grand scheme of things. Little did they know what was in store for their little insignificant world in the millennia that followed.